A vending machine for canned soda takes only 5 and 10 coins. There are five times as many 10 peso coins than 5 peso coins in the machine. The total value of the coins is 440 pesos. How many of each kind are there?
Answer by mananth(16946)

5 peso coins ------------- x
10 peso coins -------------5x
..
Value = 440
..
5x+10*5x=440
5x+50x=440
55x= 440
/55
x= 8--- 5 peso coins
5x = 5*8 = 40-- 10 peso coins
...
m.ananth@hotmail.ca